If the scores for a test have a mean of 70 and a standard deviation of 12, find the percentage of scores that will fall below 60.  

P( x < 70) = P( z < ___(j)_________) = ______(k)_____________
Answer #1

Here = 70

= 12

P(x < 60) = P[z < (60 - 70)/12] = P(z < -0.833)

here using the NORMSDIST function or using z table

P(x < 60) = P[z < (60 - 70)/12] = P(z < -0.833) = 0.2023